OpenStudy (anonymous):

solve log5(x^2-6)=log5(x+2)

OpenStudy (apoorvk):

take antilog both sides, and solve.

OpenStudy (anonymous):

is the answer -4 though

OpenStudy (anonymous):

you get x^2-x-8 = 0 since the bases are the same

OpenStudy (apoorvk):

Nope. Solve the quadratic.^

OpenStudy (anonymous):

so x = 3.72 or x = -2.37

OpenStudy (anonymous):

yes... those are solutions to the quadratic... but only one of them is a solution to the log equation.

OpenStudy (anonymous):

log5(x^2-6)=log5(x+2) since the bases are the same (x^2-6)=(x+2) simplifying : x^2-x-8 = 0 solve quadratic equation using quadratic formula : you will get either x = 3.72 or x = -2.37 since log cannot be negative : therefore : ANS : x=3.72

OpenStudy (anonymous):

Opps, sorry dude, input error in calculator X= 3.372281323 Sorry, was to quick .. So, ANS = 3.37 I'am new, interface is so cumbersome !